OpenStudy (anonymous):

Use 6 unit multipliers to convert 31 cubic inches to cubic yards

OpenStudy (ash2326):

@ckotchou74 Welcome to Open Study Let's solve this. How many inches make a foot?

OpenStudy (anonymous):

12

OpenStudy (ash2326):

Great @ckotchou74 Tell me how many foots make a yard?

OpenStudy (anonymous):

3 feet

OpenStudy (ash2326):

Great so how many inches make a yard??

OpenStudy (anonymous):

36 inches

OpenStudy (ash2326):

Ok we have 31 cubic inches or 31 inches^3 so it can be written as \[31 inch \times inch \times inch\] Do you agree with this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

how many yards make one inch?

OpenStudy (anonymous):

whtr you talking about

OpenStudy (ash2326):

\[ 36 inches = 1yard\] so \[1 inch=1/36 yards\]

OpenStudy (anonymous):

ohhhh

OpenStudy (ash2326):

@ckotchou74 so \[31 inch \times inch \times inch\] now 1 inch= 1/36 yard so \[31 \times 1/36 yard \times 1/36 yard \times 1/36 yard\] Can you solve now?

OpenStudy (anonymous):

yes thank you